### v3

# On the uniqueness of the maximum parsimony tree for data with few substitutions within the NNI neighborhood

Abstract.
Estimating species relationship trees, so-called phylogenetic trees, from aligned sequence data (such as DNA, RNA, or proteins) is one of the main aims of evolutionary biology. However, tree reconstruction criteria like maximum parsimony do not necessarily lead to unique trees and in some cases even fail to recognize the “correct” tree (i.e., the tree on which the data was generated).
On the other hand, a recent study has shown that for an alignment containing precisely those characters (sites) which require up to two substitutions on a given tree, this tree will be the unique maximum parsimony tree.
It is the aim of the present manuscript to generalize this recent result in the following sense: We show that for a tree with $n$ leaves, as long as $k<\frac{n}{8}+\frac{6}{5}-\frac{1}{10}\sqrt{\frac{5}{16}n^{2}+4}$ (or, equivalently, $n>8k-\frac{46}{5}+\frac{2}{5}\sqrt{40k-31}$), the maximum parsimony tree for the alignment containing all characters which require (up to or precisely) $k$ substitutions on a given tree $T$ will be unique in the NNI neighborhood of $T$ and it will coincide with $T$, too. In other words, within the NNI neighborhood of $T$, $T$ is the unique most parsimonious tree for said alignment. This partially answers a recently published conjecture affirmatively.
Keywords. phylogenetic tree, maximum parsimony, Buneman theorem
MSC identifiers. 05C05, 05C90, 92B05
1. Introduction
One of the main aims of mathematical phylogenetics is the reconstruction of an evolutionary relationship tree, also often called a phylogenetic tree, of a given species set $X$ based on some given data. Typically, the data are provided in the form of aligned sequence data, like DNA, RNA, proteins or binary characteristics (the latter are sometimes used to denote the absence or presence of certain morphological characteristics in species). The columns of such so-called alignments are often referred to as characters or sites.
While there are various different tree reconstruction methods available [10, 14, 2], methods based on the maximum parsimony (MP) principle are often better in settings in which the underlying evolutionary model is non-homogeneous [8]. It is also often assumed that MP is able to identify the “correct” tree (i.e., the one that has generated the data) whenever the number substitutions is relatively small [11].
The above mentioned maximum parsimony principle seeks the tree which requires as few character state changes along its edges as possible, i.e., in this sense it tries to minimize the number of mutations/substitutions needed to explain the evolution of the species set under investigation. For a given character $f$ and a given tree $T$, the minimum number of substitutions needed to explain $f$ on $T$ is often referred to as parsimony score of $f$ on $T$.
This biological context has recently inspired several mathematical publications. In particular, the findings of [6], in which for a given tree $T$ the alignment $A_{k}(T)$ consisting of all characters of parsimony score $k$ on $T$ was analyzed, led to the conjecture that whenever $n\geq 4k+1$ (or, equivalently, whenever $k\leq\frac{n-1}{4}$), $T$ is the unique maximum parsimony tree for $A_{k}(T)$, i.e., MP will recover $T$ uniquely:
Conjecture 1 (Conjecture 1 from [3]).
Let $T$ be a binary phylogenetic $X$-tree with $|X|=n$. Let $k<\frac{n}{4}$. Then, $T$ is the unique maximum parsimony tree for $A_{k}(T)$.
Conjecture 1 was published in [3] and partially proven there for the cases $k=1$ and $k=2$.
Subsequently, several publications analyzed cases for which $A_{k}(T)$ is unique for $T$. Clearly, if two trees $T$ and $T^{\prime}$ share the same $A_{k}$-alignment, i.e., if $A_{k}(T)=A_{k}(T^{\prime})$, no tree reconstruction method will be able to recover the generating tree uniquely. In [3] it was shown that the $A_{k}$-alignment is unique for $k\leq 2$ and all possible values of $n$. In [4], this result was generalized to all values of $k\geq 1$ and $n\geq 20k$. The latter result was improved by [13] to $n\geq 4k$. In [13], additionally the so-called NNI neighborhood of phylogenetic trees $T$ was analyzed, and it was shown that whenever $n\geq 2k+3$, the $A_{k}$-alignment $A_{k}(T)$ of a given tree $T$ is unique within its NNI neighborhood (i.e., small changes of the tree do not lead to trees with the same $A_{k}$-alignment). In the same publication, the authors also presented an example of two NNI neighbors which both have $22k+$ 2 leaves and share the same $A_{k}$-alignment. Moreover, already in [6] as well as in [3], an example of a tree $T$ with $n=8$ leaves and $k=2$ can be found for which, while the $A_{k}$-alignment of $T$ is unique, another tree has a smaller (and thus “better”) parsimony score than $T$, and this tree is an NNI neighbor of $T$ and thus differs only slightly from $T$, cf. Example 1.
Example 1.
Figure 1 shows a tree $T$ and one of its NNI neighbors $T^{\prime}$. It was mentioned in [6] and further analyzed in [3] that $T^{\prime}$ has a strictly smaller parsimony score for $A_{2}(T)$ than $T$.
Figure 1. Two trees, $T$ (left) and $T^{\prime}$ (right), which are NNI neighbors (which can be seen by swapping the subtree with leaves 3 and 4 with the one with leaves 7 and 8 around edge $e$ in $T$) and for which the parsimony score of $A_{2}(T)$ is smaller on $T^{\prime}$ than on $T$.
Mathematically, it is a natural question to ask if the results from [3] concerning the uniqueness of the maximum parsimony tree for $k=2$ can be generalized to $k>2$. It is the main aim of the present manuscript to show that, at least within the NNI neighborhood of a tree, as long as $k<\frac{n}{8}+\frac{6}{5}-\frac{1}{10}\sqrt{\frac{5}{16}n^{2}+4}$, this generalization indeed holds. This partially answers Conjecture 1 affirmatively.
Moreover, we will re-visit Example 1, i.e., the setting of a tree that is not a maximum parsimony tree for its own $A_{k}$-alignment as one of its NNI neighbors has a smaller parsimony score. We will show that this example can be extended in the sense that such examples exist for all $n\geq 8$ and for $k$ in the magnitude of $\frac{n}{2}$ (for exact values of $k$, see Table 1). We conclude the present manuscript by discussing our results and highlightling some interesting paths for future research.
2. Preliminaries
2.1. Definitions and basic concepts
2.1.1. Phylogenetic trees
We start with some notation. Recall that a phylogenetic tree $T=(V,E)$ on a species set $X=\{1,\ldots,n\}$ is a connected acyclic graph with vertex set $V$ and edge set $E$ whose leaves are bijectively labeled by $X$. Such a tree $T$ is also often referred to as phylogenetic $X$-tree. It is called rooted if it contains one designated root node $\rho$ and (unrooted) otherwise. Moreover, it is called binary if all its inner nodes have degree 3, except in the rooted case, in which the root must have degree 2 (except in the special case of $|X|=1$ in which the root is at the same time the only leaf of the tree). Note that we consider two phylogenetic $X$-trees $T=(V,E)$ and $T^{\prime}=(V^{\prime},E^{\prime})$ to be isomorphic, denoted $T\simeq T^{\prime}$, if there exists a map $f:V\rightarrow V^{\prime}$ such that $e=\{u,v\}\in E\Longleftrightarrow\{f(u),f(v)\}\in E^{\prime}$ and with the additional property that $f(x)=x$ for all $x\in X$. In the rooted case, we also must have $f(\rho)=\rho^{\prime}$, where $\rho$ and $\rho^{\prime}$ denote the roots of $T$ and $T^{\prime}$, respectively. In other words, $f$ is a graph isomorphism which preserves the leaf labelling, and – if applicable – the root position.
Throughout this manuscript, whenever we refer to a tree $T$, unless stated otherwise, we always mean an unrooted binary phylogenetic $X$-tree. Whenever we mean a rooted tree, we explicitly state this. However, note that sometimes it is useful to root an unrooted tree by adding an extra root node. In particular, for any (unrooted) binary phylogenetic $X$-tree $T$ we call the rooted binary tree $T_{e}$ a rooted version of $T$ with respect to $e$ if $e$ is an edge of $T$ and if $T_{e}$ can be derived from $T$ by subdividing edge $e$ by adding a new degree-2 root vertex $\rho_{e}$ to $e$. A depiction of this procedure can be found in Figure 2.
Figure 2. Subdividing an edge $e$ of a phylogenetic tree $T$ by adding a new degree-2 vertex $\rho_{e}$ turns $T$ into a rooted phylogenetic tree $T_{e}$ with root $\rho_{e}$.
Recall that a rooted tree comes with an inherent hierarchy, which implies that the vertices incident to a vertex $v$ which are not contained on the path from $v$ to the root $\rho$ of the given tree, are often referred to as children of $v$ and $v$ as their parent.
2.1.2. Characters and alignments
Now that we have characterized the objects we seek to reconstruct, namely phylogenetic trees, we need to specify the kind of data used to do so. In order to do this, recall that a character $f$ is a function from the taxon set $X$ to a set $\mathcal{C}$ of character states, i.e., $f:X\rightarrow\mathcal{C}$. Note that a finite sequence of characters is also often referred to as alignment in biology. While in most biological cases, the order of the characters in an alignment plays an important role, for our purpose it suffices to simply define an alignment as a multiset of characters. In this manuscript, we will only be concerned with binary characters (and thus also binary alignments), i.e., without loss of generality $\mathcal{C}=\{a,b\}$. Thus, a binary character $f:X\rightarrow\{a,b\}$ assigns to each leaf of the tree a corresponding state. Moreover, sometimes two alignments $A$ and $B$ are concatenated to form a new alignment $A.B$. This concatenated alignment in our context is simply the union of the two multisets $A$ and $B$.
The most important alignment we will consider, the so-called $A_{k}$-alignment $A_{k}(T)$ of a given tree $T$, will be defined at the end of Section 2.1.3.
2.1.3. Maximum parsimony and the Fitch algorithm
Now that we have the data (in the form of characters and alignments) and the objects we seek to reconstruct (namely phylogenetic trees), we need to specify an optimization criterion which we can use to achieve that. In order to do that, we first need to understand what an extension of a character is: An extension of such a character $f$ on a phylogenetic $X$-tree $T=(V,E)$ with vertex set $V$ is a map $g:V\rightarrow\{a,b\}$ such that $g(x)=f(x)$ for all $x\in X$. Thus, while $f$ only assigns states to the leaves, $g$ assigns states to all inner vertices of $T$, but it agrees with $f$ on the leaves. Moreover, we call $ch(g)=|\{\{u,v\}\in E,\,g(u)\neq g(v)\}|$ the changing number or substitution number of $g$ on $T$.
Now, the concept we need for tree reconstruction in the present manuscript is maximum parsimony (MP), which is based on the so-called parsimony score $l(f,T)$ of a character $f$ on a tree $T$. Here, $l(f,T)=\min\limits_{g}ch(g,T)$, where the minimum runs over all extensions $g$ of $f$ on $T$. The parsimony score of an alignment $A=\{f_{1},\ldots,f_{m}\}$ of characters is then defined as: $l(A,T)=\sum\limits_{i=1}^{m}l(f_{i},T)$. Moreover, a maximum parsimony tree, or MP tree for short, of an alignment $A$ is defined as $\operatorname*{arg\,min}_{T}\{l(A,T)\}$, where the minimum runs over all phylogenetic $X$-trees $T$ defined on the same set of taxa as $A$. In other words, an MP tree of an alignment $A$ is a tree minimizing the parsimony score of $A$ amongst all phylogenetic trees with the taxa of $A$ as leaves.
Note that given a binary tree $T$ and a character $f$, the parsimony score can be efficiently calculated in linear time using the well-known Fitch algorithm [5, 7]. This algorithm assigns a set of states to all inner vertices and minimizes the required number of changes. It is based on Fitch’s parsimony operation which we explain now. Therefore, let $\mathcal{C}$ be a non-empty finite set of character states and let $A,B\subseteq\mathcal{C}$. Then, Fitch’s parsimony operation *** is defined by:
$A*B\coloneqq\begin{cases}A\cap B,&\text{if }A\cap B\neq\emptyset,\\ A\cup B,&\text{otherwise.}\end{cases}$
Using this operation, the Fitch algorithm works as follows. Given a binary phylogenetic tree $T$ and a character $f$, if $T$ is not already rooted, the algorithm first adds a degree-2 root to one of the edges of $T$, cf. Figure 2. It then proceeds with the rooted tree, starting with the leaves. It assigns a set to each leaf containing precisely the state assigned to this leaf by $f$. For instance, if $f(1)=a$, leaf $1$ gets assigned set $S(f,T,1)=\{a\}$. The algorithm then proceeds as follows: In each step, it considers all vertices $v$ whose two children have already been assigned a set, say $A$ and $B$. Then, $v$ is assigned the set $S(T,f,v):=A*B$. This step is continued “upwards” (i.e., from the leaves to the root) along the tree until the root $\rho$ is assigned a set, which is denoted by $S(f,T,\rho)$. Note that throughout the present manuscript, when we want to describe the set $S$ assigned to a vertex $v$, by a slight abuse of notation, we often write $S(v)$ instead of $S(f,T,v)$ whenever there is no ambiguity. Ultimately, the parsimony score $l(f,T)$ then simply equals the number of times the Fitch operation had to use the union instead of the intersection [5]. An example for the Fitch algorithm is given by Figure 3.
Figure 3. An example of the Fitch algorithm. Here, character $f$ with $f(1)=a$, $f(2)=a$, $f(3)=b$, $f(4)=b$, $f(5)=a$, $f(6)=b$, $f(7)=b$, $f(8)=a$ is mapped onto the leaves the rooted version $T_{e}$ of $T$ from Figure 2. The Fitch algorithm then starts at the leaves and considers their state assignment as sets. It proceeds with those inner vertices whose children have already been assigned a set and applies the Fitch operation. If a union has to be taken, the counter goes up. Note that not all $\{a,b\}$-sets are the result of a union. In the figure, union nodes are marked with an asterix. As there are three such nodes, we have $l(f,T_{e})=l(f,T)=3$ and thus $f\in A_{3}(T)$.
The final and possibly most important concept we wish to introduce in this section is the following: For a given tree $T$, we define $A_{k}(T)$ to be the set consisting of all binary characters $f$ with $l(f,T)=k$. Following [3], we also refer to $A_{k}(T)$ as the alignment induced by $T$ and $k$.
2.1.4. Specific notions in the context of NNI
When considering a phylogenetic $X$-trees, it is often useful to consider small changes to the tree to see if in how far this affects, for instance, the parsimony score of certain alignments. In this manuscript, we will therefore consider nearest neighbor interchange moves or NNI moves for short. An NNI move simply takes an inner edge $e$ of a binary phylogenetic $X$-tree $T$, i.e., an edge that is not incident to a leaf of $T$, and swaps two of the four subtrees of $T$ which we get when deleting the precisely four edges adjacent to $e$ in a way that the resulting tree is not isomorphic to $T$, i.e., in a way that changes the tree. An illustration of NNI moves can be found in Figure 4. A tree resulting from $T$ by performing one NNI move is called an NNI neighbor of $T$, and all NNI neighbors of $T$ together with $T$ form the NNI neighborhood of $T$. Note that this implies that we consider $T$ to belong to its own neighborhood even though $T$ is not its own neighbor.
Figure 4. A tree $T$ and its two NNI neighbors concerning inner edge $e$: Tree $T^{\prime}$ can be derived from $T$ by swapping subtrees $T_{2}$ and $T_{4}$, whereas $T^{\ast}$ can be derived by swapping subtrees $T_{2}$ and $T_{3}$. All subtrees are schematically depicted as triangles.
It is the main aim of the present manuscript to show that if $T^{\prime}$ is an NNI neighbor of $T$, we have $l(A_{k}(T),T)<l(A_{k}(T),T^{\prime})$ if $k$ is sufficiently small (relative to the leaf number $n$ of $T$). In order to prove this, we now define certain subtrees of $T$ and $T^{\prime}$ and parameters $\delta$ and $\delta^{\prime}$ which we will frequently refer to later on.
Let $T$ and $T^{\prime}$ be phylogenetic $X$-trees such that $T$ and $T^{\prime}$ are NNI neighbors. Let $e=\{u,v\}$ be the inner edge of $T$ around which the NNI move has to be performed in order to turn $T$ into $T^{\prime}$. Then, the removal of $e$ from $T$ disconnects $T$ into two rooted subtrees. In both of these subtrees the root has degree 2 (as $e$ was an inner edge), so we can further subdivide these subtrees. Without loss of generality, we denote by $T_{1}$ and $T_{2}$ the two subtrees resulting from the deletion of $u$ and its incident edges, and similarly, we denote by $T_{3}$ and $T_{4}$ the subtrees resulting from the deletion of $v$ and its incident edges (i.e., $T$ looks like in Figure 4). The root of each $T_{i}$ is denoted by $\rho_{i}$ for $i=1,\ldots,4$.
Now let $f$ be a binary character on $X$, and let $f_{i}$ denote the restriction of $f$ on $T_{i}$ for $i=1,\ldots,4$. Then, we define function $\delta(f,T,e)$ as follows:
(1) $\delta(f,T,e):=l(f,T)-\sum\limits_{i=1}^{4}l(f_{i},T_{i}).$
Informally, $\delta(f,T,e)$ describes the number of changes of $f$ on $T$ which are not contained in any of the $T_{i}$ ($i=1,\ldots,4$) induced by $e$, but which are rather located “around $e$ itself. We will further investigate $\delta$ subsequently and also exploit its properties to derive our main result. However, as a short-hand, whenever there is no ambiguity concerning $f$ and $e$, we often use the short-hands $\delta=\delta(f,T,e)$ and $\delta^{\prime}=\delta(f,T^{\prime},e^{\prime})$, where $e^{\prime}$ denotes the unique edge in $T^{\prime}$ that is not contained in $T$ (i.e., the new edge resulting from the NNI move).
2.2. Known results
It is the main aim of the present manuscript to generalize the following two results concerning $k\leq 2$ to the case $k\geq 3$, at least within an NNI neighborhood of the given tree $T$.
The first result is the following theorem which is partially based on the famous splits equivalence theorem by Buneman [1, 10].
Theorem 1 (adapted from Corollary 1 and Theorem 3 of [3]).
Let $T$ and $T^{\prime}$ be two binary phylogenetic $X$-trees. Then, the following two statements are equivalent:
  • $T\simeq T^{\prime}$
  • $A_{1}(T)=A_{1}(T^{\prime})$
If additionally we have $|X|\geq 9$, then these statements are also equivalent to the following statement.
  • $A_{2}(T)=A_{2}(T^{\prime})$.
The second result we will generalize to the case $k\geq 3$ within the NNI neighborhood of $T$ is the following corollary, which does not only consider alignments of characters with parsimony score precisely $k$, but instead those of characters with parsimony score up to $k$.
Corollary 1 (Corollary 3 in [3]).
Let $T$ be a binary phylogenetic $X$-tree with $|X|\geq 9$. Then, $T$ is the unique maximum parsimony tree for the alignments $A_{0}.A_{1}(T)$, $A_{0}.A_{2}(T)$, $A_{1}(T).A_{2}(T)$ and $A_{0}.A_{1}(T).A_{2}(T)$.
The proof of the above lemma presented in [3] heavily relies on the following lemma, which we will also use subsequently.
Lemma 1 (Lemma 2 in [3]).
Let $T$ be a binary phylogenetic $X$-tree such that $T$ is a maximum parsimony tree of alignments $A$ and $B$ and for one of them even unique with this property. Then, $T$ is also the unique maximum parsimony tree of the concatenated alignment $A.B$.
A basic result that we need throughout this manuscript is the following theorem [^1] Theorem 2 has been adapted from its original version, most prominently by adding the final bullet point. However, it can be easily seen that this is correct as each rooted binary tree with $n$ leaves has $n-1$ inner (i.e., non-leaf) vertices. Therefore, $n-1$ is a hypothetical upper bound for the number of union vertices during the Fitch algorithm and thus also for the parsimony score. This is why a tree with $n$ leaves cannot have any characters of parsimony score more than $n-1$., which for a rooted binary tree counts the number of binary characters that have parsimony score $k$ and additionally assign the root a certain set $M$ during the Fitch algorithm.
Theorem 2 (adapted from Theorem 1 in [12]).
Let $T$ be a rooted binary phylogenetic $X$-tree with $|X|=n$. Let $k\in\mathbb{N}_{\geq 1}$. Let $N_{M}(T,k)$ denote the number of binary characters $f:X\rightarrow\{a,b\}$ on $T$ which fulfill both of the following properties:
  • $l(f,T)=k$ and
  • $S(f,T,\rho)=M$.
Then, we have:
  • If $M=\{a\}$ or $M=\{b\}$ and $k\leq n-1$, we have:
    $N_{M}(T,k)=\binom{n-k-1}{k}\cdot 2^{k}.$
  • If $M=\{a,b\}$ and $k\leq n-1$, we have:
    $N_{M}(T,k)=\binom{n-k-1}{k-1}\cdot 2^{k}.$
  • If $k>n-1$, we have $N_{M}(T,k)=0$.
Equipped with all notions and known results stated above, we finally have the tools necessary to derive new results concerning the uniqueness of the MP tree within NNI neighborhoods.
3. Results
3.1. Investigating the NNI neighborhood when $k$ is sufficiently small
It is the main aim of this manuscript to show that within its NNI neighborhood and for a suitable choice of $k$, every tree $T$ gets uniquely recovered by MP from $A_{k}(T)$. In other words, we want to show that if $T^{\prime}$ is an NNI neighbor of $T$, we have $l(A_{k}(T),T)<l(A_{k}(T),T^{\prime})$ (if $k$ is sufficiently small). This result is formally stated by the following theorem.
Theorem 3.
Let $T$ and $T^{\prime}$ be (unrooted) binary phylogenetic $X$-trees (with $|X|=n$) such that $T$ and $T^{\prime}$ are NNI neighbors. Let $k\in\mathbb{N}_{0}$ such that $k<\frac{n}{8}+\frac{6}{5}-\frac{1}{10}\sqrt{\frac{5}{16}n^{2}+4}$. Then, we have: $l(A_{k}(T),T)<l(A_{k}(T),T^{\prime})$. In other words, $T$ is the unique maximum parsimony tree of $A_{k}(T)$ within its NNI neighborhood.
Before we will prove this theorem, we first state the following observation.
Observation 1.
The upper bound for $k$ given by Theorem 3 is somewhat complicated. However, it can be easily shown (e.g., using a computer algebra system like Mathematica [9]), that $k<\frac{n}{8}+\frac{6}{5}-\frac{1}{10}\sqrt{\frac{5}{16}n^{2}+4}$ translates to $n>10k-12+2\sqrt{5k^{2}-12k+8}$. For all non-negative values of $k$, the latter term is strictly smaller than $15k$. Thus, the theorem in particular shows that if $n\geq 15k$, the MP tree of $A_{k}(T)$ is unique within the NNI neighborhood of $T$.
Next, we need to derive some tools (provided by the following lemma and proposition) which we later on require to prove Theorem 3. In particular, we first investigate the relationship between $T$ and its NNI neighbors more in-depth using $\delta$ as defined in Section 2.1.4.
Lemma 2.
Let $T$ be an unrooted binary phylogenetic $X$-tree (with $|X|=n$) with an inner edge $e=\{u,v\}$ such that $T$ is as sketched in Figure 4, i.e., $e$ splits $T_{1}$ and $T_{2}$ away from $T_{3}$ and $T_{4}$. We denote the root of $T_{i}$ with $\rho_{i}$ for $i=1,\ldots,4$. Let $f$ be a binary character on $X$ and let $S$ be the assignment of sets to the inner vertices of $T$ according to the Fitch algorithm when applied to the rooted version $T_{e}$ of $T$. Then, we have:
$\delta(f,T,e)=\begin{cases}0\mbox{ if $S$ fulfills (i) }\\ 2\mbox{ if $S$ fulfills (ii) }\\ 1\mbox{ else,}\end{cases}$
where
  1. (i)
    • $S(\rho_{1})=S(\rho_{2})=S(\rho_{3})=S(\rho_{4})$ or
    • at least three of the sets $S(\rho_{1})$, $S(\rho_{2})$, $S(\rho_{3})$, $S(\rho_{4})$ equal $\{a,b\}$ or
    • all of the sets $S(\rho_{1})$, $S(\rho_{2})$, $S(\rho_{3})$, $S(\rho_{4})$ are either contained in $\{\{a\},\{a,b\}\}$ or in $\{\{b\},\{a,b\}\}$;
  2. (ii)
    • $S(\rho_{1})=S(\rho_{4})\neq S(\rho_{2})=S(\rho_{3})$ and $|S(\rho_{i})|=1$ for all $i=1,\ldots,4$ or
    • $S(\rho_{1})=S(\rho_{3})\neq S(\rho_{2})=S(\rho_{4})$ and $|S(\rho_{i})|=1$ for all $i=1,\ldots,4$.
Proof.
  1. (i)
    • If $S(\rho_{1})=S(\rho_{2})=S(\rho_{3})=S(\rho_{4})$, then obviously $S(\rho_{e})=S(u)=S(v)=S(\rho_{1})=S(\rho_{2})=S(\rho_{3})=S(\rho_{4})$, which shows that $l(f,T)=\sum\limits_{i=1}^{4}l(f_{i},T_{i})$. Thus, by Equation (1), we have $\delta(f,T,e)=0$.
    • Similarly, if at least three of the sets $S(\rho_{1})$, $S(\rho_{2})$, $S(\rho_{3})$, $S(\rho_{4})$ equal $\{a,b\}$, then at least one of the sets $S(u)$, $S(v)$ equals $\{a,b\}$, too, and the respective other set can be obtained by an intersection, too. As $S(u)$ or $S(v)$ equals $\{a,b\}$, in any case, $S(\rho)$ can then be obtained by an intersection as well. So we have $l(f,T)=\sum\limits_{i=1}^{4}l(f_{i},T_{i})$ and thus again by Equation (1), we have $\delta(f,T,e)=0$.
    • Now let all of the sets $S(\rho_{1})$, $S(\rho_{2})$, $S(\rho_{3})$, $S(\rho_{4})$ be contained in $\{\{a\},\{a,b\}\}$ (else swap the roles of $a$ and $b$ if the case $\{\{b\},\{a,b\}\}$ shall be considered). Then, in all subsequent steps of the Fitch algorithm, either there is a pairing of $\{a\}$ and $\{a\}$ or of $\{a\}$ and $\{a,b\}$ or of $\{a,b\}$ and $\{a,b\}$. In all of these cases, the intersection will be taken, which again shows $l(f,T)=\sum\limits_{i=1}^{4}l(f_{i},T_{i})$ and thus again $\delta(f,T,e)=0$.
  2. (ii)
    Now let $S(\rho_{1})=S(\rho_{4})\neq S(\rho_{2})=S(\rho_{3})$ and $|S(\rho_{i})|=1$ for all $i=1,\ldots,4$ (note that the second subcase with the roles of $\rho_{3}$ and $\rho_{4}$ interchanged works analogously). Without loss of generality, we have $\{a\}=S(\rho_{1})=S(\rho_{4})$ and $\{b\}=S(\rho_{2})=S(\rho_{3})$ (else swap the roles of $a$ and $b$). Obviously, the Fitch algorithm will then assign $S(u)=\{a,b\}$ and $S(v)=\{a,b\}$, and both are unions ($S(\rho_{e})=\{a,b\}$, on the other hand, is only an intersection and thus does not count). Therefore, we derive $l(f,T)=\sum\limits_{i=1}^{4}l(f_{i},T_{i})+2$ (as we had to add two unions to the partial parsimony scores) and thus, again by Equation (1), we have $\delta(f,T,e)=2$.
  3. (iii)
    Now we consider all other cases:
    • If two of the sets $S(\rho_{1})$, $S(\rho_{2})$, $S(\rho_{3})$, $S(\rho_{4})$ are equal to $\{a\}$ and the other two sets are $\{b\}$, but we are not in case (ii), i.e., if we have $S(\rho_{1})=S(\rho_{2})\neq S(\rho_{3})=S(\rho_{4})$, then the Fitch algorithm assigns $S(u)=S(\rho_{1})=S(\rho_{2})$ as well as $S(v)=S(\rho_{3})=S(\rho_{4})$ by intersection. Subsequently, we derive $S(\rho_{e})=\{a,b\}$ by taking the union. Thus, we would have $l(f,T)=\sum\limits_{i=1}^{4}l(f_{i},T_{i})+1$ and therefore $\delta(f,T,e)=1$.
    • If two of the sets $S(\rho_{1})$, $S(\rho_{2})$, $S(\rho_{3})$, $S(\rho_{4})$ are equal to $\{a,b\}$ and the other two sets are $\{a\}$ and $\{b\}$, respectively, there are two possibilities: Either both $\{a,b\}$ sets’ corresponding vertices are adjacent to the same node (say, $u$) or not. If they are, this node will be an intersection and it will also be assigned $\{a,b\}$ by the Fitch algorithm. The other two sets, however, would then also be adjacent to the same vertex, say $v$, and they are $\{a\}$ and $\{b\}$. So clearly, the Fitch algorithm would assign their union $\{a,b\}$ to their adjacent vertex, say $v$, so $S(v)=\{a,b\}$. For $S(\rho_{e})$, however, the intersection could be taken again.
      On the other hand, if the two $\{a,b\}$ sets’ corresponding vertices are not adjacent to the same node, then one of them has a $\{a\}$ node as a sibling and one of them has a $\{b\}$ node as a sibling. Thus, we would get either $S(u)=\{a\}$ and $S(v)=\{b\}$ or vice versa, but this would not require any union taking. $S(\rho_{e})$, however, would equal $\{a,b\}$ and it would be the result of taking a union.
      So in both cases, we would have to take the union once, so we have $l(f,T)=\sum\limits_{i=1}^{4}l(f_{i},T_{i})+1$ (as we have to add one union to the partial parsimony scores) and thus, again by Equation (1), we have $\delta(f,T,e)=1$.
    • If precisely one of the sets $S(\rho_{1})$, $S(\rho_{2})$, $S(\rho_{3})$, $S(\rho_{4})$ is equal to $\{a,b\}$ and two of the sets are equal to $\{a\}$ and one is equal to $\{b\}$ (or vice versa), then whichever of the sets is a sibling to the $\{a,b\}$ set will use an intersection, so the parent of the $\{a,b\}$ gets assigned a set of cardinality 1, and it equals the sibling of the $\{a,b\}$ set.
      The other pairing, however, might consist of two equal sets of cardinality 1, in which case their parent gets assigned this very set by intersection, too. As this set differs from the set assigned to the parent of $\{a,b\}$, the Fitch algorithm requires a union to determine $S(\rho_{e})=\{a,b\}$.
      Otherwise, the other pairing must consist of a pairing of two differing sets $\{a\}$ and $\{b\}$, so their parent will get assigned $\{a,b\}$ by taking the union. Consequently, in the next step, $S(\rho_{e})$ will be assigned the singleton set that occurred twice amongst $S(\rho_{1})$, $S(\rho_{2})$, $S(\rho_{3})$, $S(\rho_{4})$, and this would happen by intersection.
      Thus, in both of these cases, we have $l(f,T)=\sum\limits_{i=1}^{4}l(f_{i},T_{i})+1$ (as we had to add one union to the partial parsimony scores) and thus we have $\delta(f,T,e)=1$. This completes the proof.
While Lemma 2 only considered $\delta$ in the context of one tree, the following proposition will derive some crucial consequences concerning $\delta$ when considering a tree and one of its NNI neighbors together. This proposition will be the main ingredient in the proof of Theorem 3.
Proposition 1.
Let $T$, $e$ and $T_{1},\ldots T_{4}$ be like in Lemma 2. Let $T^{\prime}$ be an NNI neighbor of $T$ derived by swapping two of the subtrees around $e$ (which generates an edge $e^{\prime}$ in $T^{\prime}$ corresponding to $e$ in $T$, cf. Figure 4). Let $f$ be a binary character on $X$. Then, we have: If $l(f,T)\neq l(f,T^{\prime})$, then either $\delta(f,T,e)=2$ and $\delta(f,T^{\prime},e^{\prime})=1$ or vice versa. In particular, we must have that two of the sets $S(\rho_{1})$, $S(\rho_{2})$, $S(\rho_{3})$, $S(\rho_{4})$ are equal to $\{a\}$ and the other two are equal to $\{b\}$.
Proof.
Let $T$ and $T^{\prime}$ be two NNI neighbors as described in the statement of the proposition; without loss of generality, $T^{\prime}$ can be derived from $T$ by swapping $T_{2}$ and $T_{4}$. Let $f$ be a binary character such that $l(f,T)\neq l(f,T^{\prime})$.
Now, if the conditions of (i) in Lemma 2 are met, both $\delta(f,T,e)$ and $\delta(f,T^{\prime},e^{\prime})$ equal 0, so in particular, these values coincide. However, in this case, by Equation (1), we have $0=\delta(f,T,e)=l(f,T)-\sum\limits_{i=1}^{4}l(f_{i},T_{i})$ as well as $0=\delta(f,T^{\prime},e^{\prime})=l(f,T^{\prime})-\sum\limits_{i=1}^{4}l(f_{i},T_{i}^{\prime})$. Using $T_{i}=T_{i}^{\prime}$ for all $i=1,\ldots,4$ (as $T$ and $T^{\prime}$ are NNI neighbors), this implies
$l(f,T)=\sum\limits_{i=1}^{4}l(f_{i},T_{i})=\sum\limits_{i=1}^{4}l(f_{i},T_{i}^{\prime})=l(f,T^{\prime}),$
which contradicts the assumption $l(f,T)\neq l(f,T^{\prime})$. Thus this case need not be considered.
If, however, neither conditions (i) nor (ii) of Lemma 2 are met, we have three possibilities (cf. Part (iii) of the proof of Lemma 2):
  • We have $S(\rho_{1})=S(\rho_{2})\neq S(\rho_{3})=S(\rho_{4})$ and $|S(\rho_{i})|=1$ for all $i=1\ldots,4$, or
  • two of the sets $S(\rho_{1})$, $S(\rho_{2})$, $S(\rho_{3})$, $S(\rho_{4})$ are equal to $\{a,b\}$ and the other two sets are $\{a\}$ and $\{b\}$, respectively, or
  • precisely one of the sets $S(\rho_{1})$, $S(\rho_{2})$, $S(\rho_{3})$, $S(\rho_{4})$ is equal to $\{a,b\}$ and two of the sets are equal to $\{a\}$ and one is equal to $\{b\}$ (or vice versa).
In both latter cases, by Lemma 2, however, we have $\delta(f,T,e)=\delta(f,T^{\prime},e^{\prime})=1$, so again both values are equal. Note that this is due to the fact that these conditions on $S(\rho_{i})$ ($i=1,\ldots,4$) are independent of the positions of subtrees $T_{i}$ ($i=1,\ldots,4$) in the tree. However, if $\delta(f,T,e)=\delta(f,T^{\prime},e^{\prime})$, as above this would imply $l(f,T)=l(f,T^{\prime})$, a contradiction to our assumption. So these cases need not be considered further.
In the first of the above three cases, however, by Lemma 2, we have $\delta(f,T,e)=1$ and $\delta(f,T^{\prime},e^{\prime})=2$ (as the roles of $\rho_{2}$ and $\rho_{4}$ are reversed in $T^{\prime}$). Therefore, in this case, the two values indeed differ, and indeed we have that either $S(\rho_{1})=S(\rho_{2})=\{a\}$ and $S(\rho_{3})=S(\rho_{4})=\{b\}$ or the roles of $a$ and $b$ are interchanged. Moreover, while we have $\delta(f,T,e)=1$, we have $\delta(f,T^{\prime},e^{\prime})=2$ in this case (as $T_{2}$ and $T_{4}$ as well as $T_{1}$ and $T_{3}$ are grouped by $e^{\prime}$, respectively, and both require a union to form $\{a,b\}$ from $\{a\}$ and $\{b\}$, respectively. So in this case, the statement of the proposition is true and there remains nothing more to show.
So the only remaining cases in which we can have $\delta(f,T,e)\neq\delta(f,T^{\prime},e^{\prime})$ are the cases induced by (ii) of Lemma 2: Here, we must have
  • $S(\rho_{1})=S(\rho_{4})\neq S(\rho_{2})=S(\rho_{3})$ and $|S(\rho_{i})|=1$ for all $i=1,\ldots,4$ or
  • $S(\rho_{1})=S(\rho_{3})\neq S(\rho_{2})=S(\rho_{4})$ and $|S(\rho_{i})|=1$ for all $i=1,\ldots,4$.
This already proves the second assertion of the proposition.
To prove the first part, remember that $T_{2}$ and $T_{4}$ where swapped to derive $T^{\prime}$ from $T$.
Now if the second of the above cases held, i.e., if we had $S(\rho_{1})=S(\rho_{3})\neq S(\rho_{2})=S(\rho_{4})$ and $|S(\rho_{i})|=1$ for all $i=1,\ldots,4$, then $\delta(f,T,e)=2$, but $\delta(f,T^{\prime},e^{\prime})=2$, too. This is due to the fact that we can apply Lemma 2 to $T^{\prime}$ and $e^{\prime}$, too, but the roles of $T_{2}$ and $T_{4}$ are reversed (compared to $T$). Thus, both values of $\delta$ are equal. As $\delta(f,T,e)=\delta(f,T^{\prime},e^{\prime})$ implies $l(f,T)=l(f,T^{\prime})$ as explained above, this contradicts our assumption and thus cannot be the case.
So instead, we must have $S(\rho_{1})=S(\rho_{4})\neq S(\rho_{2})=S(\rho_{3})$ and $|S(\rho_{i})|=1$ for all $i=1,\ldots,4$. Applying Lemma 2 to $T$ shows that $\delta(f,T,e)=2$, and applying Lemma 2 to $T^{\prime}$ shows that $\delta(f,T,e)=1$ in this case. Therefore, in this case, the two values indeed differ.
So in all cases in which $\delta(f,T,e)$ and $\delta(f,T^{\prime},e^{\prime})$ differ, we have that one of them equals 1 and the other value equals 2, and in all of these cases, two of the sets $S(\rho_{1})$, $S(\rho_{2})$, $S(\rho_{3})$, $S(\rho_{4})$ are equal to $\{a\}$ and the other two sets are equal to $\{b\}$. This completes the proof. ∎
We are now finally in the position to prove Theorem 3.
Proof of Theorem 3.
Let $k\in\mathbb{N}_{0}$ such that $k<\frac{n}{8}+\frac{6}{5}-\frac{1}{10}\sqrt{\frac{5}{16}n^{2}+4}$. If $k\leq 2$, according to Theorem 1, there remains nothing to show. Thus, we may assume in the following that $k\geq 3$.
Now, let $T$ be a binary phylogenetic $X$-tree with $|X|=n$ and let $e=\{u,v\}$ be an inner edge of $T$ inducing subtrees $T_{1}$ and $T_{2}$ adjacent to $u$ and $T_{3}$ and $T_{4}$ adjacent to $v$ when $e$ is removed. In the following, we denote the roots of $T_{i}$ by $\rho_{i}$, respectively.
Let $T^{\prime}$ be the NNI neighbor of $T$ resulting from swapping $T_{2}$ and $T_{4}$ around $e$, leading to an inner edge $e^{\prime}$ in $T^{\prime}$. We need to show that $l(A_{k}(T),T)<l(A_{k}(T),T^{\prime})$, or, in other words, we want to show that $l(A_{k}(T),T)-l(A_{k}(T),T^{\prime})<0$.
Our proof strategy is now as follows: We will analyze $A_{k}(T)$ and count characters of certain categories. Note that $A_{k}(T)$ may contain various characters that have the same parsimony score on $T$ and $T^{\prime}$. These characters do not contribute to the difference of the parsimony scores and thus need not be considered as they cancel out in the above mentioned difference. So we will only consider characters $f\in A_{k}(T)$ for which we have $k=l(f,T)\neq l(f,T^{\prime})$. By Proposition 1, these characters fall into precisely two categories:
  1. (1)
    Either $\delta(f,T,e)=2$ and $\delta(f,T^{\prime},e^{\prime})=1$ or
  2. (2)
    $\delta(f,T,e)=1$ and $\delta(f,T^{\prime},e^{\prime})=2$.
In both cases, again by Proposition 1 and its proof, we know that in both cases, two of the sets $S(\rho_{1})$, $S(\rho_{2})$, $S(\rho_{3})$ and $S(\rho_{4})$ equal $\{a\}$ and the other two equal $\{b\}$. Moreover, the characters in Category (1) have $S(\rho_{1})\neq S(\rho_{2})$ as well as $S(\rho_{3})\neq S(\rho_{4})$, whereas the characters in Category (2) have $S(\rho_{1})=S(\rho_{2})$ as well as $S(\rho_{3})=S(\rho_{4})$.
So our proof strategy is to count the number $c_{1}$ of characters of Category (1) and the number $c_{2}$ of characters of Category (2) and show that $c_{2}>c_{1}$. This will lead to the following conclusion:
(2) $l(A_{k}(T),T)-l(A_{k}(T),T^{\prime})$ $=(c_{1}\cdot k+c_{2}\cdot k)-(c_{1}\cdot(k-1)+c_{2}\cdot(k+1))$
$=(c_{1}\cdot k+c_{2}\cdot k)-(c_{1}\cdot k+c_{2}\cdot k-c_{1}+c_{2})$
$=c_{1}-c_{2}$
$<0.$
Here, the first equality is due to the fact that – as explained above – characters which do not fall into Categories (1) or (2) need not be considered in the difference. Moreover, this first equality exploits the fact that for characters in Category (1), we have $\delta(f,T^{\prime},e^{\prime})=1$ and thus, using Definition 1 and the fact that $\sum\limits_{i=1}^{4}l(f_{i},T_{i})=\sum\limits_{i=1}^{4}l(f_{i},T_{i}^{\prime})$, $l(f,T^{\prime})=k-1$ for all $f$ in $A_{k}(T)$ which belong to Category (1), i.e., they benefit from the NNI move from $T$ to $T^{\prime}$ by a decrease in their parsimony score from $k$ to $k-1$. Similarly, for characters in Category (2), we have $l(f,T^{\prime})=k+1$, i.e., such characters suffer from the NNI move from $T$ to $T^{\prime}$ by an increase in their parsimony score from $k$ to $k+1$.
So all that remains to be shown is that indeed we have $c_{1}<c_{2}$.
In order to calculate $c_{1}$, in the light of Proposition 1, as explained above we need to count the number of characters with $S(\rho_{1})\neq S(\rho_{2})$ and $S(\rho_{3})\neq S(\rho_{4})$ and such that for all $i=1,\ldots,4$ we have $|S(\rho_{i})|=1$. More precisely, this means that we have four possibilities:
  1. (i)
    $S(\rho_{1})=\{a\}$, $S(\rho_{2})=\{b\}$, $S(\rho_{3})=\{a\}$, $S(\rho_{4})=\{b\}$, or
  2. (ii)
    $S(\rho_{1})=\{a\}$, $S(\rho_{2})=\{b\}$, $S(\rho_{3})=\{b\}$, $S(\rho_{4})=\{a\}$, or
  3. (iii)
    $S(\rho_{1})=\{b\}$, $S(\rho_{2})=\{a\}$, $S(\rho_{3})=\{a\}$, $S(\rho_{4})=\{b\}$, or
  4. (iv)
    $S(\rho_{1})=\{b\}$, $S(\rho_{2})=\{a\}$, $S(\rho_{3})=\{b\}$, $S(\rho_{4})=\{a\}$.
Note that in each of these possibilities, we need to sum up over all possible values of $l(f_{i},T_{i})$. More precisely, we know that $l(f,T)=k$ (as $f\in A_{k}(T)$), and we know that $\delta(f,T,e)=2$. So by Equation (1), we must have $\sum\limits_{i=1}^{4}l(f_{i},T_{i})=k-2$. However, these $k-2$ changes can be distributed amongst the four subtrees $T_{1},\ldots,T_{4}$ in many possible ways, and we have to consider them all. To facilitate notation, in the following we let $k_{i}:=l(f_{i},T_{i})$ for $i=1,\ldots,4$. Then we must have $k_{1}+k_{2}+k_{3}+k_{4}=k-2$. Moreover, if we denote by $n_{i}$ the number of leaves in $T_{i}$ (for $i=1,\ldots,4$), then we must have $n_{1}+n_{2}+n_{3}+n_{4}=n$.
Furthermore, note that the number of characters with $l(f_{i},T_{i})=k_{i}$ is independent of the fact if $S(\rho_{i})=\{a\}$ or $S(\rho_{i})=\{b\}$. By Theorem 2, it simply equals $N_{a}(T_{i},k_{i}):=N_{\{a\}}(T_{i},k_{i})=\binom{n_{i}-k_{i}-1}{k_{1}}\cdot 2^{k_{i}}$.
Thus, in summary, we derive:
(6) $c_{1}$ $=\sum\limits_{\begin{subarray}{c}(k_{1},k_{2},k_{3},k_{4}):\\ k_{1}+k_{2}+k_{3}+k_{4}=k-2\\ k_{i}\in\mathbb{N}_{0}\end{subarray}}4\cdot N_{a}(n_{1},k_{1})\cdot N_{a}(n_{2},k_{2})\cdot N_{a}(n_{3},k_{3})\cdot N_{a}(n_{4},k_{4}).$
Now in order to calculate $c_{2}$, again in the light of Proposition 1, we need to count the number of characters with $S(\rho_{1})=S(\rho_{2})\neq S(\rho_{3})=S(\rho_{4})$ and such that for all $i=1,\ldots,4$ we have $|S(\rho_{i})|=1$. More precisely, this means that we have two possibilities:
  1. (i)
    $S(\rho_{1})=\{a\}$, $S(\rho_{2})=\{a\}$, $S(\rho_{3})=\{b\}$, $S(\rho_{4})=\{b\}$, or
  2. (ii)
    $S(\rho_{1})=\{b\}$, $S(\rho_{2})=\{b\}$, $S(\rho_{3})=\{a\}$, $S(\rho_{4})=\{a\}$.
Note that in each of these possibilities, we need to sum up over all possible values of $l(f_{i},T_{i})$. More precisely, we know that $l(f,T)=k$ (as $f\in A_{k}(T)$), and we know that $\delta(f,T,e)=1$. So by Equation (1), we must have $\sum\limits_{i=1}^{4}l(f_{i},T_{i})=k-1$. However, as above, these $k-1$ changes can be distributed amongst the four subtrees $T_{1},\ldots,T_{4}$ in many possible ways, and we have to consider them all. However, as this time, we have $\delta(f,T,e)=1$, one of the $k_{i}$ needs to get incrememented by 1 (compared to the $c_{1}$ case).
Thus, in summary, we derive:
(10) $c_{2}=\sum\limits_{\begin{subarray}{c}(k_{1},k_{2},k_{3},k_{4}):\\ k_{1}+k_{2}+k_{3}+k_{4}=k-2\\ k_{i}\in\mathbb{N}_{0}\end{subarray}}2\cdot$ $\left(N_{a}(n_{1},k_{1}+1)\cdot N_{a}(n_{2},k_{2})\cdot N_{a}(n_{3},k_{3})\cdot N_{a}(n_{4},k_{4})\right.$
$+N_{a}(n_{1},k_{1})\cdot N_{a}(n_{2},k_{2}+1)\cdot N_{a}(n_{3},k_{3})\cdot N_{a}(n_{4},k_{4})$
$+N_{a}(n_{1},k_{1})\cdot N_{a}(n_{2},k_{2})\cdot N_{a}(n_{3},k_{3}+1)\cdot N_{a}(n_{4},k_{4})$
$\left.+N_{a}(n_{1},k_{1})\cdot N_{a}(n_{2},k_{2})\cdot N_{a}(n_{3},k_{3})\cdot N_{a}(n_{4},k_{4}+1)\right).$
We will now show $c_{2}>c_{1}$ by actually showing something stronger: We will show that each of the summands in the sum of $c_{1}$ is strictly smaller than each of the summands in the sum of $c_{2}$. This will complete the proof.
Note that we have by Theorem 2:
  • $N_{a}(n_{i},k_{i})=\binom{n_{i}-k_{i}-1}{k_{i}}\cdot 2^{k_{i}}=\frac{(n_{i}-k_{i}-1)!}{(n_{i}-2k_{i}-1)!\cdot k!}\cdot 2^{k_{i}},$
  • $N_{a}(n_{i},k_{i}+1)=\binom{n_{i}-(k_{i}+1)-1}{k_{i}+1}\cdot 2^{k_{i}+1}=\binom{n_{i}-k_{i}-2}{k_{i}+1}\cdot 2^{k_{i}+1}=\frac{(n_{i}-k_{i}-2)!}{(n_{i}-2k_{i}-3)!\cdot(k+1)!}\cdot 2^{k_{i}+1}$.
Thus, as long as $n_{i}\neq k_{i}+1$ (so that the denominator does not go to 0), we get:
(11) $N_{a}(n_{i},k_{i}+1)$ $=2\cdot N_{a}(n_{i},k_{i})\cdot\frac{(n_{i}-2k_{i}-2)\cdot(n_{i}-2k_{i}-1)}{(n_{i}-k_{i}-1)\cdot(k_{i}+1)}.$
Moreover, we have $N_{a}(n_{i},k_{i})\geq 0$ and $N_{a}(n_{i},k_{i}+1)\geq 0$ for all $n_{i}\in\mathbb{N}_{\geq 1}$, $k_{i}\in\mathbb{N}_{0}$ by Theorem 2.
Last but not least, by the pigeonhole principle, we know that at least one of the $T_{i}$ has $n_{i}\geq\left\lceil\frac{n}{4}\right\rceil$ many leaves (because we distribute $n$ leaves of $T$ among the four subtrees $T_{1},\ldots,T_{4}$). Without loss of generality, we therefore may assume that $n_{1}\geq\left\lceil\frac{n}{4}\right\rceil$. While we do not now how big $k_{1}$ is, we at least know that $k_{1}\leq k-2<n_{1}-1$:
[…]
(12) $=\left(\frac{n_{1}}{2}-\frac{4}{5}-\frac{1}{10}\sqrt{5n_{1}^{2}+4}\right)$
$\leq n_{1}-1$
Here, the second and third inequalities hold as $\frac{n}{4}\leq\left\lceil\frac{n}{4}\right\rceil\leq n_{1}$ and as the function $f(n)=\frac{1}{2}\cdot\frac{n}{4}+\frac{6}{5}-\frac{1}{10}\sqrt{5\cdot\left(\frac{n}{4}\right)^{2}+4}$ is monotonically increasing. The last inequality is true as it holds for all $n_{1}\geq 1$ (and we know that $n_{i}\geq 1$ for all $i=1,\ldots,4$ as each $T_{i}$ contains at least one leaf; otherwise $e$ would not be an inner edge).
So in particular, we have $n_{1}\neq k_{1}+1$, which means we can use Equation (11) to derive $N_{a}(n_{1},k_{1}+1)$ from $N_{a}(n_{1},k_{1})$.
Let now $(k_{1},k_{2},k_{3},k_{4})$ be a tuple such that $k_{1}+k_{2}+k_{3}+k_{4}=k-2$ and $k_{i}\in\mathbb{N}_{0}$ for all $i=1,\ldots,4$. Let $c_{i}^{(k_{1},k_{2},k_{3},k_{4})}$ denote the summand of $c_{i}$ corresponding to the tuple $(k_{1},k_{2},k_{3},k_{4})$ for $i=1,2$. Then, we have:
$c_{2}^{(k_{1},k_{2},k_{3},k_{4})}\geq$ $2\cdot N_{a}(n_{1},k_{1}+1)\cdot N_{a}(n_{2},k_{2})\cdot N_{a}(n_{3},k_{3})\cdot N_{a}(n_{4},k_{4})$
=(11)${=}2\cdot\left(2\cdot N_{a}(n_{1},k_{1})\cdot\frac{(n_{1}-2k_{1}-2)\cdot(n_{1}-2k_{1}-1)}{(n_{1}-k_{1}-1)\cdot(k_{1}+1)}\right)\cdot N_{a}(n_{2},k_{2})\cdot N_{a}(n_{3},k_{3})\cdot N_{a}(n_{4},k_{4})$ )
$=\frac{(n_{1}-2k_{1}-2)\cdot(n_{1}-2k_{1}-1)}{(n_{1}-k_{1}-1)\cdot(k_{1}+1)}\cdot 4\cdot N_{a}(n_{1},k_{1})\cdot N_{a}(n_{2},k_{2})\cdot N_{a}(n_{3},k_{3})\cdot N_{a}(n_{4},k_{4})$
$=\frac{(n_{1}-2k_{1}-2)\cdot(n_{1}-2k_{1}-1)}{(n_{1}-k_{1}-1)\cdot(k_{1}+1)}\cdot c_{1}^{(k_{1},k_{2},k_{3},k_{4})}.$
Thus, we definitely have $c_{2}^{(k_{1},k_{2},k_{3},k_{4})}>c_{1}^{(k_{1},k_{2},k_{3},k_{4})}$ if $\frac{(n_{1}-2k_{1}-2)\cdot(n_{1}-2k_{1}-1)}{(n_{1}-k_{1}-1)\cdot(k_{1}+1)}>1$, or, in other words, if $(n_{1}-2k_{1}-2)\cdot(n_{1}-2k_{1}-1)>(n_{1}-k_{1}-1)\cdot(k_{1}+1)$. Using a computer algebra system like Mathematica [9], one can easily see that this does indeed hold in case $k_{1}<\frac{n_{1}}{2}-\frac{4}{5}-\frac{1}{10}\sqrt{5n_{1}^{2}+4}$. However, we have already seen in Equation (12) that this inequality holds.
This shows that $c_{2}^{(k_{1},k_{2},k_{3},k_{4})}>c_{1}^{(k_{1},k_{2},k_{3},k_{4})}$ for all tuples $(k_{1},k_{2},k_{3},k_{4})$ with $k_{i}\in\mathbb{N}_{0}$ for all $i=1,\ldots,4$ such that $k_{1}+k_{2}+k_{3}+k_{4}=k-2$. This in turn shows that $c_{2}>c_{1}$, which in the light of Equation (2) proves that $l(A_{k}(T),T)>l(A_{k}(T),T^{\prime})$ and thus completes the proof. ∎
We conclude this section with the following result, generalizing Theorem 3 from alignments $A_{k}(T)$, which contain all characters with parsimony score precisely $k$ on $T$, to alignments of characters of score at most $k$ on $T$.
Corollary 2.
Let $k\in\mathbb{N}$ Let $T$ be a binary phylogenetic $X$-tree with $|X|=n>10k-12+2\sqrt{5k^{2}-12k+8}$. Let $T^{\prime}$ be an NNI-neighbor of $T$. For $i=0,\ldots,k$, let $m_{i}\in\mathbb{N}_{0}$ and let $A$ be an alignment consisting of concatenations of $A_{i}(T)$ such that $A_{i}(T)$ is contained $m_{i}$ times in $A$ (for $i=0,\ldots,k$). Then, $l(f,T)<l(f,T^{\prime})$.
Proof.
The corollary is a direct consequence of Observation 1 as well as Theorems 1 (observing that if $k=2$ and $|X|=n>10k-12+2\sqrt{5k^{2}-12k+8}$, we have $n>12$, which in particular ensures that $n\geq 9$ as required by Theorem 1 is not a problem) and 3 together with Lemma 1. ∎
3.2. Investigating the NNI neighborhood when $k$ is “too large”
In this section, we want to use our findings from the previous section to derive values for $k$ for which $T$ is not the most parsimonious tree for $A_{k}(T)$ in its own NNI neighborhood. In order to do so, we will exploit Equation (2). In particular, we will show that whenever $k$ assumes a certain value (depending on $n$), we can construct a tree $T$ which has an NNI neighbor $T^{\prime}$ such that $l(A_{k}(T),T)>l(A_{k}(T),T^{\prime})$. In other words, in these settings, $T$ is not even contained in the set of most parsimonious trees for its own $A_{k}$-alignment. We now state this formally in the following theorem, which will generalize Example 1, in which we had $n=8$ and $k=2=\frac{n-4}{2}$.
Theorem 4.
Let $n\in\mathbb{N}_{\geq 8}$ and let $k$ be chosen according to Table 1. Then, there exists a binary phylogenetic $X$-tree $T$ with $|X|=n$ such that $T$ has an NNI neighbor $T^{\prime}$ with $l(A_{k}(T),T)>l(A_{k}(T),T^{\prime})$. In other words, $T^{\prime}$ is more parsimonious for $A_{k}(T)$ than $T$.
condition on $n$ choice of $k$
$(n\mod 8)\equiv 0$ $\frac{n-4}{2}$
$(n\mod 8)\in\{1,7\}$ $\frac{n-3}{2}$
$(n\mod 8)\in\{2,6\}$ $\frac{n-2}{2}$
$(n\mod 8)\in\{3,5\}$ $\frac{n-1}{2}$
$n\mod 8\equiv 4$ $\frac{n}{2}$
Table 1. Choices of $k$ for different values of $n$ to generate a “bad case” as described in Theorem 4.
Before we can prove this theorem, we need another technical lemma.
Lemma 3.
Let $n$, $k\in\mathbb{N}_{0}$. Let
$g(n,k)=\begin{cases}\binom{n-k-1}{k}&\mbox{if $k\leq n-1$}\\ 0&\mbox{else.}\end{cases}$
Then, if $g(n,k)>0$, we have $k<\left\lceil\frac{n}{2}\right\rceil-\frac{1}{2}$.
Proof.
In the following, let $g(n,k)>0$.
We first consider the case that $n$ is even. In this case, we have $\left\lceil\frac{n}{2}\right\rceil-\frac{1}{2}=\frac{n}{2}-\frac{1}{2}=\frac{n-1}{2}\not\in\mathbb{N}_{0}$, which means that the lemma requires $k\leq\frac{n-2}{2}$.
Assume this is not the case, i.e., assume $k\geq\frac{n}{2}$ (but we still have $g(n,k)>0$ and thus must have $k\leq n-1$ by definition of $g$).
So we consider the case $\frac{n}{2}\leq k\leq n-1$. In this case, we have $0\leq n-k-1\leq\frac{n}{2}-1<k$ and thus $g(n,k)=\binom{n-k-1}{k}=0$ (as the upper value of the binomial coefficient is non-negative but smaller than the lower one). This is a contradiction to the assumption $g(n,k)>0$, which proves the lemma for the case in which $n$ is even.
Next, we consider the case that $n$ is odd. In this case, we have $\left\lceil\frac{n}{2}\right\rceil-\frac{1}{2}=\frac{n+1}{2}-\frac{1}{2}=\frac{n}{2}\not\in\mathbb{N}_{0}$, which means that the lemma requires $k\leq\frac{n-1}{2}$.
Assume this is not the case, i.e., assume $k\geq\frac{n+1}{2}$ (but we still have $g(n,k)>0$ and thus must have $k\leq n-1$ by definition of $g$).
So we consider the case $\frac{n+1}{2}\leq k\leq n-1$. In this case, we have $0\leq n-k-1\leq\frac{n-1}{2}-1<k$ and thus $g(n,k)=\binom{n-k-1}{k}=0$ (as the upper value of the binomial coefficient is non-negative but smaller than the lower one). This is a contradiction to the assumption $g(n,k)>0$, which completes the proof. ∎
We are now finally in a position to prove Theorem 4.
Proof of Theorem 4.
We explicitly construct a tree $T$ and its NNI neighbor $T^{\prime}$ with the described properties. We do this by using a strategy closely related to that of the proof of Theorem 3.
Let $n=8m+b$ with $m\in\mathbb{N}_{\geq 1}$ and $b\in\{0,\ldots,7\}$. We first choose four parameters $n_{1}$, $n_{2}$, $n_{3}$ and $n_{4}$ as follows:
$n_{1}:=\begin{cases}2m&\mbox{if $b=0$}\\ 2m+1&\mbox{if $b\in\{1,\ldots,4\}$}\\ 2m+2&\mbox{if $b\geq 5$},\end{cases}$ $,\hskip 14.22636ptn_{2}:=\begin{cases}2m&\mbox{if $b\leq 1$}\\ 2m+1&\mbox{if $b\in\{2,\ldots,5\}$}\\ 2m+2&\mbox{if $b\geq 6$},\end{cases}$
$n_{3}:=\begin{cases}2m&\mbox{if $b\leq 2$}\\ 2m+1&\mbox{if $b\in\{3,\ldots,6\}$}\\ 2m+2&\mbox{if $b=7$},\end{cases}$ $,\hskip 14.22636ptn_{4}:=\begin{cases}2m&\mbox{if $b\leq 3$}\\ 2m+1&\mbox{if $b\geq 4$}.\end{cases}$
Note that in all cases, we have $n_{1}+n_{2}+n_{3}+n_{4}=n$, which is due to the choice of these parameters:
  • If $b=0$ and $n=8m+0$, we have $n_{1}=n_{2}=n_{3}=n_{4}=2m$.
  • If $b=1$ and $n=8m+1$, we have $n_{1}=2m+1$ and $n_{2}=n_{3}=n_{4}=2m$.
  • If $b=2$ and $n=8m+2$, we have $n_{1}=n_{2}=2m+1$ and $n_{3}=n_{4}=2m$.
  • If $b=3$ and $n=8m+3$, we have $n_{1}=n_{2}=n_{3}=2m+1$ and $n_{4}=2m$.
  • If $b=4$ and $n=8m+4$, we have $n_{1}=n_{2}=n_{3}=n_{4}=2m+1$.
  • If $b=5$ and $n=8m+5$, we have $n_{1}=2m+2$ and $n_{2}=n_{3}=n_{4}=2m+1$.
  • If $b=6$ and $n=8m+6$, we have $n_{1}=n_{2}=2m+2$ and $n_{3}=n_{4}=2m+1$.
  • If $b=7$ and $n=8m+7$, we have $n_{1}=n_{2}=n_{3}=2m+2$ and $n_{4}=2m+1$.
To construct $T$ and $T^{\prime}$, we now start with a single edge $e$, around which we attach rooted binary subtrees $T_{1}$, $T_{2}$, $T_{3}$ and $T_{4}$ with $n_{1}$, $n_{2}$, $n_{3}$ and $n_{4}$ leaves, respectively. We make sure that in $T$, $T_{1}$ and $T_{2}$ are together on the same side of $e$, whereas $T_{3}$ and $T_{4}$ are on the other side of $e$. We then arbitrarily label the leaves such that $X=\{1,\ldots,n\}$. This is possible as $n_{1}+n_{2}+n_{3}+n_{4}=n$. We now define $T^{\prime}$ as the tree resulting from $T$ by swapping $T_{2}$ and $T_{4}$.
Now we choose $k$ according to Table 1. In particular, this leads to:
  • If $b=0$ and $n=8m+0$, we have $k=\frac{n-4}{2}=\frac{8m-4}{2}=4m-2$.
  • If $b=1$ and $n=8m+1$, we have $k=\frac{n-3}{2}=\frac{(8m+1)-3}{2}=4m-1$.
  • If $b=2$ and $n=8m+2$, we have $k=\frac{n-2}{2}=\frac{(8m+2)-2}{2}=4m$.
  • If $b=3$ and $n=8m+3$, we have $k=\frac{n-1}{2}=\frac{(8m+3)-1}{2}=4m+1$.
  • In all cases in which $b\geq 4$, we have $k=4m+2$.
Moreover, the characters in Category (1) have $S(\rho_{1})\neq S(\rho_{2})$ as well as $S(\rho_{3})\neq S(\rho_{4})$, whereas the characters in Category (2) have $S(\rho_{1})=S(\rho_{2})$ as well as $S(\rho_{3})=S(\rho_{4})$.
From now on, we follow the strategy of the proof of Theorem 3. In particular, we observe again that characters in $A_{k}(T)$ which have the same parsimony score on $T$ and $T^{\prime}$ do not contribute to the difference $l(A_{k}(T),T)-l(A_{k}(T),T^{\prime})$. We have seen in said proof that only the $c_{1}$ many characters of Category 1 (two of the sets $S(\rho_{1})$, $S(\rho_{2})$, $S(\rho_{3})$ and $S(\rho_{4})$ equal $\{a\}$ and the other two equal $\{b\}$ and $S(\rho_{1})\neq S(\rho_{2})$ as well as $S(\rho_{3})\neq S(\rho_{4})$) and the $c_{2}$ many characters of Category 2 (two of the sets $S(\rho_{1})$, $S(\rho_{2})$, $S(\rho_{3})$ and $S(\rho_{4})$ equal $\{a\}$ and the other two equal $\{b\}$ and $S(\rho_{1})=S(\rho_{2})$ as well as $S(\rho_{3})=S(\rho_{4})$) contribute to this difference. In particular, we know $l(A_{k}(T),T)-l(A_{k}(T),T^{\prime})=c_{1}-c_{2}$.
We now want to calculate $c_{1}$ and $c_{2}$. We will start with showing that $c_{1}>0$.
In the following, we consider the rooted version $T_{e}$ of $T$ and denote by $\rho_{i}$ the root vertex of $T_{i}$ for $i\in\{a,b,c,d\}$. We know by Proposition 1 that for all characters $f\in A_{k}(T)$ from Category (1), we need to make sure that $|S(\rho_{i})|=1$ for all $i=1,\ldots,4$ according to the Fitch algorithm. Combining Theorem 2 with Lemma 3, we know that in order for this to be possible, the number $k_{i}$ (with $i\in\{a,b,c,d\}$) of union nodes assigned by the Fitch algorithm within $T_{i}$ cannot exceed $\left\lceil\frac{n_{i}}{2}\right\rceil-\frac{1}{2}$. In our cases, this translates to:
  • If $b=0$, $n=8m+0$ and $k=4m-2$, we have $n_{1}=n_{2}=n_{3}=n_{4}=2m$ and thus $k_{1},k_{2},k_{3},k_{4}\leq m-1$. Note that this implies $\sum\limits_{i=1}^{4}k_{i}\leq k-2$.
  • If $b=1$, $n=8m+1$ and $k=4m-1$, we have $n_{1}=2m+1$ and $n_{2}=n_{3}=n_{4}=2m$, and thus $k_{1}\leq m$ and $k_{2},k_{3},k_{4}\leq m-1$. Note that this implies $\sum\limits_{i=1}^{4}k_{i}\leq k-2$.
  • If $b=2$, $n=8m+2$ and $k=4m$, we have $n_{1}=n_{2}=2m+1$ and $n_{3}=n_{4}=2m$, and thus $k_{1},k_{2}\leq m$ and $k_{3},k_{4}\leq m-1$. Note that this implies $\sum\limits_{i=1}^{4}k_{i}\leq k-2$.
  • If $b=3$, $n=8m+3$ and $k=4m+1$, we have $n_{1}=n_{2}=n_{3}=2m+1$ and $n_{4}=2m$, and thus $k_{1},k_{2},k_{3}\leq m$ and $k_{4}\leq m-1$. Note that this implies $\sum\limits_{i=1}^{4}k_{i}\leq k-2$.
  • In all cases with $b\geq 4$, we have $k=4m+2$ and $k_{1},k_{2},k_{3},k_{4}\leq m$. Note that this implies $\sum\limits_{i=1}^{4}k_{i}\leq k-2$.
Now, in order for $f$ to be contained in $A_{k}(T)$, we require:
$k=l(f,T)=\underbrace{\sum\limits_{i=1}^{4}k_{i}}_{\leq k-2}+\underbrace{\delta(f,T,e)}_{\leq 2}.$
This is only possible if all the $k_{i}$ reach their respective upper bound (cf. above bullet points) and $\delta(f,T,e)=2$. This is due to the fact that for each such $f$, we have $k=l(f,T)=\sum\limits_{i=1}^{4}l(f_{i},T_{i})+\delta(f,T,e)$.
We now finally turn our attention to $c_{1}$, of which we want to show that it is strictly positive. By Equation (6) as well as the above considerations (namely that there is only one choice for each $k_{i}$, namely its upper bound according to the above bullet point list, making the summation in Equation (6) redundant), we know that
$c_{1}=4\cdot N_{a}(n_{1},k_{1})\cdot N_{a}(n_{2},k_{2})\cdot N_{a}(n_{3},k_{3})\cdot N_{a}(n_{4},k_{4}).$
As the $k_{i}$ were all chosen to equal the maximum possible value not exceeding $\left\lceil\frac{n_{i}}{2}\right\rceil-\frac{1}{2}$, by Theorem 2 and Lemma 3 we know that $N_{a}(n_{i},k_{i})>0$ for all $i$. This immediately shows that $c_{1}>0$.
Next, we will show that $c_{2}=0$. By Equation (10) as well as the observation from above that we only have one choice for each $k_{i}$ (namely its respective upper bound) in order to reach $k_{1}+k_{2}+k_{3}+k_{4}=k-2$ (again making the summation in (10) redundant), we have:
$c_{2}=2\cdot$ $\left(\underbrace{N_{a}(n_{1},k_{1}+1)}_{=0}\cdot N_{a}(n_{2},k_{2})\cdot N_{a}(n_{3},k_{3})\cdot N_{a}(n_{4},k_{4})\right.$
$+N_{a}(n_{1},k_{1})\cdot\underbrace{N_{a}(n_{2},k_{2}+1)}_{=0}\cdot N_{a}(n_{3},k_{3})\cdot N_{a}(n_{4},k_{4})$
$+N_{a}(n_{1},k_{1})\cdot N_{a}(n_{2},k_{2})\cdot\underbrace{N_{a}(n_{3},k_{3}+1)}_{=0}\cdot N_{a}(n_{4},k_{4})$
$\left.+N_{a}(n_{1},k_{1})\cdot N_{a}(n_{2},k_{2})\cdot N_{a}(n_{3},k_{3})\cdot\underbrace{N_{a}(n_{4},k_{4}+1)}_{=0}\right)$
$=0.$
Here, the fact that $N_{a}(n_{i},k_{i}+1)=0$ stems from the $k_{i}$ being chosen as the maximum possible value not exceeding $\left\lceil\frac{n_{i}}{2}\right\rceil-\frac{1}{2}$. Thus, $k_{i}+1$ exceeds this value, which by Theorem 2 and Lemma 3 implies $N_{a}(n_{i},k_{i}+1)=0$.
So in summary, we have seen $c_{1}>0$ and $c_{2}=0$. As in the proof of Theorem 3, we can then calculate the difference of parsimony scores of $A_{k}(T)$ on $T$ and $T^{\prime}$ as follows:
$l(A_{k}(T),T)-l(A_{k}(T),T^{\prime})=c_{1}-c_{2}=c_{1}>0,$
which shows that $l(A_{k}(T),T^{\prime})<l(A_{k}(T),T)$. This completes the proof. ∎
4. Discussion and outlook
In this manuscript, we have shown that as long as $k<\frac{n}{8}+\frac{6}{5}-\frac{1}{10}\sqrt{\frac{5}{16}n^{2}+4}$, any binary phylogenetic tree $T$ is the unique maximum parsimony tree for $A_{k}(T)$ within its NNI neighborhood. The most obvious question for future research arising from this result is if this also holds outside of this neighborhood.
Another question is if $k$ really needs to be so small. In the proof of Theorem 3, we made $k$ so small that not only $c_{1}<c_{2}$ as required, but that each and every summand of $c_{1}$ is smaller than every summand of $c_{2}$. Relaxing this strict condition might lead to an improved bound for $k$. We conjecture that improving $k$ is indeed possible, particularly in the light of Conjecture 1 as well as the findings presented in [6].
Our second result, namely the extension of Example 1 to arbitrary values of $n\geq 8$ in Theorem 4, in which $k$ is in the magnitude of $\frac{n}{2}$ (cf. Table 1), is also of particular interest. In our construction of NNI neighbors $T$ and $T^{\prime}$ such that $l(A_{k}(T),T^{\prime})<l(A_{k}(T),T)$, in the case where $n$ is a multiple of 8 and $k=\frac{n-4}{2}$, we have $n=2k+4$. Similarly, if $n\mod 8\equiv 1$, we have $k=\frac{n-3}{2}$ and thus $n=2k+3$. However, it was shown in [13] that for all values of $n\geq 2k+3$, the $A_{k}$-alignment $A_{k}(T)$ of $T$ is unique within the NNI neighborhood of $T$. Yet our result shows that despite this uniqueness, this does not need to imply optimality. So the question whether or not $T$ is the unique maximum parsimony tree of $A_{k}(T)$ does not only depend on the question if it shares this alignment with another tree. Even in cases in which equality of scores can be excluded, surprisingly another tree might still be better than $T$ concerning the parsimony criterion – even in the NNI neighborhood of $T$. This shows that the manuscripts [4] and [13], which deal with the uniqueness of the $A_{k}$-alignment, indeed have a different flavor than [3] as well as the present manuscript, which deal with the reconstructability of $T$ from $A_{k}(T)$ with parsimony.
Last but not least, we hope that future research will close the gap between the case where $k$ is in the magnitude of $\frac{n}{2}$ and in which we know by Theorem 4 that the NNI neighborhood of $T$ can contain trees with a smaller parsimony score than $T$, and the case where $k$ is in the magnitude of $\frac{n}{15}$, in which we know by Theorem 3 and Observation 1 that this cannot happen.
Acknowledgements
MF wants to thank Mike Steel for bringing Theorem 2 to her attention, which helped to improve an earlier bound. Moreover, MF wishes to thank Mirko Wilde, Sophie Kersting, Linda Knüver and Stephan Dominique Andres for helpful discussions on this topic.
Conflict of interest
The author herewith certifies that she has no affiliations with or involvement in any organization or entity with any financial (such as honoraria; educational grants; participation in speakers’ bureaus; membership, employment, consultancies, stock ownership, or other equity interest; and expert testimony or patent-licensing arrangements) or non-financial (such as personal or professional relationships, affiliations, knowledge or beliefs) interest in the subject matter discussed in this manuscript.
Data availability statement
Data sharing is not applicable to this article as no new data were created or analyzed in this study.